Balerio Corporation's relevant range of activity is 6,000 units to 10,000 units. When it produces and sells 8,000 units, its average costs per unit are as follows:

Average Cost per Unit
Direct materials $ 6.40
Direct labor $ 3.00
Variable manufacturing overhead $ 1.30
Fixed manufacturing overhead $ 13.20
Fixed selling expense $ 2.20
Fixed administrative expense $ 1.60
Sales commissions $ 0.60
Variable administrative expense $ 0.50

Required:

a. For financial reporting purposes, what is the total amount of product costs incurred to make 8,000 units? (Do not round intermediate calculations.)

b. If 9,000 units are sold, what is the variable cost per unit sold? (Round "Per unit" answer to 2 decimal places.)

c. If 9,000 units are sold, what is the total amount of variable costs related to the units sold? (Do not round intermediate calculations. Round "Per unit" answer to 2 decimal places.)

d. If the selling price is $18.00 per unit, what is the contribution margin per unit sold? (Round "Per unit" answer to 2 decimal places.)

e. What incremental manufacturing cost will the company incur if it increases production from 8,000 to 8,001 units? (Round "Per unit" answer to 2 decimal places.)
Answer #1
a
Direct materials 51200 =8000*6.4
Direct labor 24000 =8000*3
Variable manufacturing overhead 10400 =8000*1.3
Fixed manufacturing overhead 105600 =8000*13.2
Total product cost 191200
b
Direct materials 6.40
Direct labor 3.00
Variable manufacturing overhead 1.30
Sales commissions 0.60
Variable administrative expense 0.50
Variable cost per unit sold 11.80
c
Units sold 9000
X Variable cost per unit sold 11.80
Total variable costs 106200
d
Selling price per unit 18.00
Less: Variable cost per unit sold 11.80
Contribution margin per unit 6.20
e
Direct materials 6.40
Direct labor 3.00
Variable manufacturing overhead 1.30
Incremental manufacturing cost 10.70